OpenStudy (anonymous):

integral of ln^2 (3x) / x dx???

OpenStudy (lgbasallote):

let u = ln (3x) du = 1/x dx \[\int u^2 du\]

OpenStudy (anonymous):

Thanks @lgbasallote I got the answer. :) \[1/3\ln ^{3}3x\]

OpenStudy (lgbasallote):

yup sounds right :)